Week of 200921

WLCG Operations Call details

  • For remote participation we use the Vidyo system. Instructions can be found here.

General Information

  • The purpose of the meeting is:
    • to report significant operational issues (i.e. issues which can or did degrade experiment or site operations) which are ongoing or were resolved after the previous meeting;
    • to announce or schedule interventions at Tier-1 sites;
    • to inform about recent or upcoming changes in the experiment activities or systems having a visible impact on sites;
    • to provide important news about the middleware;
    • to communicate any other information considered interesting for WLCG operations.
  • The meeting should run from 15:00 Geneva time until 15:20, exceptionally to 15:30.
  • The SCOD rota for the next few weeks is at ScodRota
  • Whenever a particular topic needs to be discussed at the operations meeting requiring information from sites or experiments, it is highly recommended to announce it by email to wlcg-scod@cernSPAMNOTNOSPAMPLEASE.ch to allow the SCOD to make sure that the relevant parties have the time to collect the required information, or invite the right people at the meeting.

Best practices for scheduled downtimes

Monday

Attendance:

  • remote: Kate (chair, DB), Julia (WLCG), Borja (monitoring), Eric C (storage), Pavlo Svirin (ATLAS), David B (IN2P3), Ben (computing), Andrew (TRIUMF), Darren (RAL), Xavier (KIT), Xin (BNL), Dave M. (FNAL)

Experiments round table:

  • ALICE -
    • NTR

  • LHCb reports ( raw view) -
    • Activity:
      • MC and WG productions; User jobs.
    • Issues:
      • RAL: Issue with one user unable to access data due to auth issue (GGUS:148701)

Sites / Services round table:

  • ASGC: nc
  • BNL: NTR
  • CNAF: Scheduled Downtime on October 1st from 07:30PM to 10:00PM CEST for an intervention to the border router, GOCDB Link
  • EGI: nc
  • FNAL: NTR
  • IN2P3: Tomorrow, site will be in downtime. CEs and SEs will be out of service all morning.
  • JINR: NTR
  • KISTI: NTR
  • KIT:
    • Last week's issues with duplicated LHCb jobs were traced back to trouble with ARC-CE-4 and have been resolved by its restart.
    • Announced site downtime October 6th/7th in GOC-DB #29457. We'll try our best to finish everything on the first day, so the second day is meant to be a buffer in case something goes wrong.
    • ARC-CEs 4 and 6 enter draining mode tomorrow and will be shut down next week GOC-DB:#29447. The remaining ARC-CEs will be decommissioned at the site downtime in October GOC-DB:#29466.
Julia has asked for the revalidation of August data, as ALICE storage disk consumption is very low.
  • NDGF: nc
  • NL-T1: NTR; unable to dial in
  • NRC-KI: nc
  • OSG: nc
  • PIC: For the upcoming Monday 28th of September, PIC will be in a complete scheduled downtime due to the following interventions: UPS intervention, Library firmware update, dCache server hardware migration. PIC will be in OUTAGE from 28-09-2020 07:00 (PIC local time) until 28-09-2020 16:00 (PIC local time). HTCondor will be stopped at the beginning of the SD.
  • RAL: NTR
  • TRIUMF: NTR

  • CERN computing services:
    • OTG0059138 VOMS Upgrade Tue Sep 22 09:30. No impact foreseen.
    • OTG0059140 Closure of remaining SLC6 Batch capacity. Jobs accepted till 24/11/2020 in line with removal of lxplus6. Capacity will be removed on 30/11/2020
  • CERN storage services:
  • CERN databases: Firewall rules will be re-deployed on the ATLAS Online Database (ATONR) hosts, sessions will need to reconnect on Friday, Sept 25th, 10.00 (OTG:0059149)
  • GGUS: NTR
  • Monitoring:
    • SAM3 dashboards stopped today as agreed in the last WLCG ops/coord meeting
    • New FTS transfers dashboard fully based on ES and using CRIC as topology (fresh data)
      • Once experiments can confirm the CRIC topology is working as expected for this use case we will proceed with new ones
      • Tentative next one: the freshly migrated SiteMon
  • MW Officer: nc
  • Networks: NTR
  • Security: nc

AOB:

Edit | Attach | Watch | Print version | History: r25 < r24 < r23 < r22 < r21 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r25 - 2020-09-22 - SangUnAhn
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2021 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
or Ideas, requests, problems regarding TWiki? use Discourse or Send feedback